# Forward KL Regularized Preference Optimization for Aligning Diffusion Policies

### v1

@@ Introduction
In solving sequential decision-making problems, Reinforcement Learning (RL) algorithms typically adopt Gaussian or deterministic policy classes in policy optimization. Although such a policy class has successfully solved various challenging tasks (Fujimoto, Meger, and Precup 2019; Lange, Gabel, and Riedmiller 2012), it can be limited in learning multi-modal policies and lead to sub-optimal behaviors in complex environments (Janner et al. 2022). Recently, diffusion models (Ho, Jain, and Abbeel 2020; Rombach et al. 2022) demonstrated superior performance compared to previous policy classes, especially for offline RL where the dataset is collected by a mixture of policies (Fu et al. 2020; Qin et al. 2022), or the embodied manipulation tasks where the policy should imitate diverse human behaviors (Chi et al. 2023; Pearce et al. 2023). Leveraging the highly expressive model capabilities, the diffusion policy exhibits strong capabilities in modeling complex behaviors.
A central problem in learning a diffusion policy is aligning the policy output with human intents (Shin, Dragan, and Brown 2023; Yuan et al. 2024). Existing methods can be roughly divided into three categories. (i) Several methods collect expert trajectories and directly learn a diffusion policy from the expert dataset (Reuss et al. 2023; Ze et al. 2024), while they usually require a large number of expert trajectories to model human behaviors. To ensure the diffusion policy generates actions with desired properties, (ii) several methods adopt conditional generation by employing the cumulative return as the condition in modeling the trajectory (Yuan et al. 2023; Ajay et al. 2023), and actions with high returns can be generated via guided sampling (Ho and Salimans 2021; Nichol and Dhariwal 2021). (iii) Other works consider diffusion policy as an actor in an RL framework and perform policy improvement by maximizing cumulative rewards. Then actions can be sampled with high $Q$-values (Wang, Hunt, and Zhou 2023; Kang et al. 2023). Although the latter two methods can learn policies from an arbitrary dataset, they rely on manually defined reward functions for each task, which can be difficult to obtain in embodied tasks (Dong et al. 2024; Gu et al. 2023). Motivated by recent advances in Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022) that align human preference with Large Language Models (LLMs) (OpenAI 2023; Touvron et al. 2023), we wonder whether the diffusion policy can be aligned with human intent directly from a collection of preference data, which removes the potentially incorrect assumption that the reward function alone drives human preferences (Knox et al. 2024; Lambert et al. 2024) and aims to guide the diffusion policy from preferences directly.
@@
Figure 1: Overal scheme of FKPD.
@@
Figure 2: Samples generated by a toy diffusion model.
@@
Next we conduct an in-depth analysis of why forward KL regularization outperforms reverse KL regularization for the aligning of diffusion policy. To this end, we track three variables in the process of alignment. The first two variables are average D-MSE of winning segments and losing segments, abbreviated as $E_{\rm winning}$ and $E_{\rm losing}$, respectively. They are capable of measuring the variation in the likelihood of winning segments and losing segments during the alignment process of diffusion policy. The third variable is implicit accuracy, which is written as
@@ Conclusions
In this paper, we propose FKPD, a method to align a diffusion policy with preference data. To tackle the OOD issue during the alignment phase, we employ the forward KL regularization, which demonstrates superior performance compared to the reverse KL regularization in the experiments. FKPD achieves the sate-of-the-art performance on open MetaWorld and D4RL locomotion tasks, indicating that it is an efficient preference aligning method. Currently, we have only validated the effectiveness of FKPD on single-modal and single-task scenarios. In future work, we aim to extend FKPD to encompass multi-task and multi-modal data.

### v2

@@ Introduction
In solving sequential decision-making problems, Reinforcement Learning (RL) algorithms typically adopt Gaussian or deterministic policy classes in policy optimization. Although such a policy class has successfully solved various challenging tasks (Fujimoto, Meger, and Precup 2019; Lange, Gabel, and Riedmiller 2012), it can be limited in learning multi-modal policies and lead to sub-optimal behaviors in complex environments (Janner et al. 2022). Recently, diffusion models (Ho, Jain, and Abbeel 2020; He et al. 2024) demonstrated superior performance compared to previous policy classes, especially for offline RL where the dataset is collected by a mixture of policies (Fu et al. 2020; Qin et al. 2022; Yuan et al. 2024a; Fan et al. 2024), or the embodied manipulation tasks where the policy should imitate diverse human behaviors (Chi et al. 2023; Pearce et al. 2023). Leveraging the highly expressive model capabilities, the diffusion policy exhibits strong capabilities in modeling complex behaviors.
A central problem in learning a diffusion policy is aligning the policy output with human intents (Shin, Dragan, and Brown 2023; Yuan et al. 2024b; Yu et al. 2024). Existing methods can be roughly divided into three categories. (i) Several methods collect expert trajectories and directly learn a diffusion policy from the expert dataset (Reuss et al. 2023; Ze et al. 2024), while they usually require a large number of expert trajectories to model human behaviors. To ensure the diffusion policy generates actions with desired properties, (ii) several methods adopt conditional generation by employing the cumulative return as the condition in modeling the trajectory (Yuan et al. 2023; Ajay et al. 2023), and actions with high returns can be generated via guided sampling (Ho and Salimans 2021; Nichol and Dhariwal 2021). (iii) Other works consider diffusion policy as an actor in an RL framework and perform policy improvement by maximizing cumulative rewards. Then actions can be sampled with high $Q$-values (Wang, Hunt, and Zhou 2023; Kang et al. 2023). Although the latter two methods can learn policies from an arbitrary dataset, they rely on manually defined reward functions for each task, which can be difficult to obtain in embodied tasks (Dong et al. 2024; Gu et al. 2023). Motivated by recent advances in Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022) that align human preference with Large Language Models (LLMs) (OpenAI 2023; Touvron et al. 2023), we wonder whether the diffusion policy can be aligned with human intent directly from a collection of preference data, which removes the potentially incorrect assumption that the reward function alone drives human preferences (Knox et al. 2024; Lambert et al. 2024) and aims to guide the diffusion policy from preferences directly.
@@
A reverse chain, constructed as $p_{\theta}(\mathbb{x}_{0:T}):=\mathcal{N}(\mathbb{x}_{T};\mathbb{0},\mathbb{I})\prod_{t=1}^{T}p_{\theta}(\mathbb{x}_{t-1}|\mathbb{x}_{t})$, is then optimized by maximizing the evidence lower bound (ELBO) defined as $\mathbb{E}_{q}\left[\ln\frac{p_{\theta}(\mathbb{x}_{0})}{q(\mathbb{x}_{1:T}|\mathbb{x}_{0})}\right]$. After training, sampling from the diffusion model consists of sampling from $\mathbb{x}_{T}$ and running the reverse diffusion chain from $t=T$ to $t=0$. Diffusion models can be extended to conditional generative models by conditioning $p_{\theta}(\mathbb{x}_{t-1}|\mathbb{x}_{t},c)$. Since there are two different types of timesteps in this work, one for the diffusion process and one for the MDP, we use superscripts to denote trajectory timesteps and subscripts to denote diffusion timesteps. We denote the diffusion-base policy via the reverse process of a conditional diffusion model as
@@
Figure 1: Overall scheme of FKPD: FKPD comprises two phases: the behavior cloning phase and the alignment phase. In the behavior clone phase, FKPD fits a diffusion policy to a preference-free dataset $\mathcal{D}$. Subsequently, in the alignment phase, FKPD employs a direct preference optimization method to align the diffusion policy with a preference dataset $\mathcal{D}_{\rm pref}$. Throughout this process, the diffusion policy is required to maintain a forward KL distance with respect to the distribution of $\mathcal{D}_{\rm pref}$
@@
Figure 2: Samples generated by a toy diffusion model. The left sub-figure displays samples generated by the initial model. The middle and right sub-figures show samples generated by the aligned models with forward and reverse KL regularization, respectively.
@@
Next we conduct an in-depth analysis of why forward KL regularization outperforms reverse KL regularization for the aligning of diffusion policy. To this end, we track three variables in the process of alignment. The first two variables are average D-MSE of winning segments and losing segments, abbreviated as $E_{\rm winning}$ and $E_{\rm losing}$, respectively.
They are capable of measuring the variation in the likelihood of winning segments and losing segments during the alignment process of diffusion policy. The third variable is implicit accuracy, which is written as
@@ Conclusions
In this paper, we propose FKPD, a method to align a diffusion policy with preference data. To tackle the OOD issue during the alignment phase, we employ the forward KL regularization, which demonstrates superior performance compared to the reverse KL regularization in the experiments. FKPD achieves the sate-of-the-art performance on open MetaWorld and D4RL locomotion tasks, indicating that it is an efficient preference aligning method. Currently, we have only validated the effectiveness of FKPD for single-modal observations, that is, the joint states of embodiment, and single-task scenarios. In future work, we aim to extend FKPD to encompass multi-task and multi-modal data. Moreover, other efficient regularization methods, such as those proposed in (Huang et al. 2024; Cen et al. 2024), have been introduced for aligning LLMs. In our future work, we plan to analyze the performance of these new regularization methods for aligning diffusion policies.
Acknowledgments
This study is supported by National Natural Science Foundation of China (Grant No.62306242). Zhao extends gratitude to professor Chenjia for providing in-depth guidance on this work. Thanks are also extended to Chenyou and the other authors for their outstanding contributions to the research.